OpenStudy (anonymous):

what is the value of X A.2 square root 3 B.3 C.3 square root 2 D.6

OpenStudy (anonymous):



OpenStudy (neer2890):

use Pythagoras theorem \[base ^{2}+height ^{2}= hypotenuse ^{2}\] so x^2=3^2+3^2=18 and x=\[3\sqrt{2}\]

OpenStudy (neer2890):

keep in mind this theorem will only be applicable in right angled triangle.

OpenStudy (anonymous):

thank u